The power of D. E. Knuth’s TEX has made typesetting of the highest quality available to the ordinary computer user, PC or otherwise. The one drawback is that the actual printing of a document can be slow. In most cases one makes many versions before the final one: it would be handy to be able to see the TEX output quickly, even in a rough form. That is the purpose of CDVI. If one has an IBM PC (or compatible) with a graphics screen (Hercules, CGA, EGA, VGA, MCGA or Olivetti/ATT) and the DVI file, then CDVI gives a screen image in seconds. CDVIO is for Olivetti/ATT 640-400 graphics, CDVI-2 is for VGA or MCGA graphics; the other are named according to the type of graphics. The restrictions are that only the Plain TEX fonts may be used and that the page dimensions are not too big.

The basic problem in writing a program to screen TEX is that the number of dots in Hercules graphics (720 by 348) is about 1/10 that of an ordinary sheet of paper with a not too fancy printer, while the CGA (640 by 200) has about 1/20 the resolution. This limitation in resolution can be overcome to some extent by vertical scrolling. In principle, horizontal scrolling could reduce the limitation even more; but in practice, legibility is considerably reduced. Nevertheless, it is possible to get about 6 inches of TEX at 10 points in the 640 horizontal dots of CGA graphics with reasonable legibility. This lacks the elegance of TEX as produced on a 300 pixel per inch device, but a first draft need not be quite so refined.

These versions of CDVI incorporate 16 fonts into the program: one only needs the program and the DVI file. The fonts occupy about 40 kb. They were made by Metafont from the CM.MF files. Best results will be obtained with versions of TEX which use the CM fonts. For generality, the program also accepts DVI files calling for AM fonts, with in many cases satisfactory results.

The versions of CDVI up to and including version 1.2 may be copied and passed on to others. CDVI 2.0 has more fonts and horizontal scrolling, as well as greater speed and convenience. Most important, however, CDVI 2.0 implements fast and sophisticated font substitution which allows a very broad range of fonts to be represented without requiring megabytes of disk space. Though it is not free, CDVI 2.0 is available at a reasonable price.

The restriction to Plain TEX fonts for the versions up to 1.2 is severe, but there is something to be said for a certain amount of discipline among those who use TEX. Not only does TEX provide an excellent means for arranging technical text on paper, but has the potential to be the best means of rapid communication of technical ideas and expressions. The proliferation of fonts hinders this second goal. Commercial interests are already trying to impose their fonts on TEX. There is no need that the beauty of TEX should interfere with its affinity for rapid communication, provided that a system of standards is accepted. The individual could choose any available fonts for hardcopy output, but for communication by electronic means he should also present a version using a standard set of fonts at specified magnifications. A database of standard DVI files would be accessible to the user of a modest PC. A file could be seen in seconds and printed in minutes.

Press the Rtn key.

TEX FONTS AVAILABLE IN CDVI

1. cmr10

000  ΓΔΘΛΞΠΣΥ  ΦΨΩfffiflffiffl  ıȷ` ´ ˇ ˘  ¯  ˚  ¸ ßæœøÆŒØ
032 !”#$%&’  ()*+,-./  01234567  89:;¡=¿?
064  @ABCDEFG  HIJKLMNO  PQRSTUVW  XYZ[“]ˆ ˙
096  ‘abcdefg  hijklmno  pqrstuvw  xyz–—˝ ˜ ¨

2. cmr7

000  ΓΔΘΛΞΠΣΥ  ΦΨΩfffiflffiffl  ıȷ` ´ ˇ ˘  ¯  ˚  ¸ ßæœøÆŒØ
032 !”#$%&’  ()*+,-./  01234567  89:;¡=¿?
064  @ABCDEFG  HIJKLMNO  PQRSTUVW  XYZ[“]ˆ ˙
096  ‘abcdefg  hijklmno  pqrstuvw  xyz–—˝ ˜ ¨

3. cmr5

000  ΓΔΘΛΞΠΣΥ  ΦΨΩfffiflffiffl  ıȷ` ´ ˇ  ˘  ¯  ˚  ¸ ßæœøÆŒØ
032 !”#$%&’  ()*+,-./  01234567  89:;¡=¿?
064  @ABCDEFG  HIJKLMNO  PQRSTUVW  XYZ[“]ˆ ˙
096  ‘abcdefg  hijklmno  pqrstuvw  xyz–—˝ ˜ ¨

4. cmsl10

000  ΓΔΘΛΞΠΣΥ  ΦΨΩfffiflffiffl  ıȷ` ´ ˇ ˘  ¯  ˚  ¸ ßæœøÆŒØ
032 !”#$%&’  ()*+,-./  01234567  89:;¡=¿?
064  @ABCDEFG  HIJKLMNO  PQRSTUVW  XYZ[“]ˆ ˙
096  ‘abcdefg  hijklmno  pqrstuvw  xyz–—˝ ˜ ¨

5. cmtt10

000  ΓΔΘΛΞΠΣΥ    ΦΨΩ↑↓'¡¿    ıȷ` ´  ˇ   ˘    ¯     ˚          ¸ ßæœøÆŒØ
032  ␣!"#$%&'    ()*+,-./    01234567    89:;<=>?
064  @ABCDEFG    HIJKLMNO    PQRSTUVW    XYZ[\]^_
096  `abcdefg    hijklmno    pqrstuvw    xyz{|}~¨

6. cmmi10

000  ΓΔΘΛΞΠΣΥ  ΦΨΩαβγδϵ  ζηθικλμν  ξπρστυϕχ
032  ψωεϑϖϱςφ  ↼↽⇀⇁ ▷◁  01234567  89.,</>⋆
064  ∂ABCDEFG  HIJKLMNO  PQRSTUVW  XYZ♭♮♯⌣⌢
096  ℓabcdefg  hijklmno  pqrstuvw  xyzıȷ℘

7. cmmi7

000  ΓΔΘΛΞΠΣΥ  ΦΨΩαβγδϵ  ζηθικλμν  ξπρστυϕχ
032  ψωεϑϖϱςφ  ↼↽⇀⇁ ▷◁  01234567  89.,</>⋆
064  ∂ABCDEFG  HIJKLMNO  PQRSTUVW  XYZ♭♮♯⌣⌢
096  ℓabcdefg  hijklmno  pqrstuvw  xyzıȷ℘

8. cmmi5

000  ΓΔΘΛΞΠΣΥ  ΦΨΩαβγδϵ  ζηθικλμν  ξπρστυϕχ
032  ψωεϑϖϱςφ  ↼↽⇀⇁ ▷◁  01234567  89.,</>⋆
064  ∂ABCDEFG  HIJKLMNO  PQRSTUVW  XYZ♭♮♯⌣⌢
096  ℓabcdefg  hijklmno  pqrstuvw  xyzıȷ℘

9. cmsy10

000  −·×∗÷⋄±∓  ⊕⊖⊗⊘⊙◯∘•  ≍≡⊆⊇≤≥⪯⪰  ∼≈⊂⊃≪≫≺≻
032  ←→↑↓↔↗↘≃  ⇐⇒⇑⇓⇔↖↙∝  ′∞∈∋△▽  ∀̸∃¬∅ℜℑ⊤⊥
064  ℵABCDEFG  HIJKLMNO  PQRSTUVW  XYZ∪∩⊎∧∨
096  ⊢⊣⌊⌋⌈⌉{}  ⟨⟩|∥↕⇕\≀  √⨿∇∫⊔⊓⊑⊒  §†‡¶♣♢♡♠

10. cmsy7

000  −·×∗÷⋄±∓  ⊕⊖⊗⊘⊙◯∘•  ≍≡⊆⊇≤≥⪯⪰  ∼≈⊂⊃≪≫≺≻
032  ←→↑↓↔↗↘≃  ⇐⇒⇑⇓⇔↖↙∝  ′∞∈∋△▽  ∀̸∃¬∅ℜℑ⊤⊥
064  ℵABCDEFG  HIJKLMNO  PQRSTUVW  XYZ∪∩⊎∧∨
096  ⊢⊣⌊⌋⌈⌉{}  ⟨⟩|∥↕⇕\≀  √⨿∇∫⊔⊓⊑⊒  §†‡¶♣♢♡♠

11. cmsy5

000  −·×∗÷⋄±∓  ⊕⊖⊗⊘⊙◯∘•  ≍≡⊆⊇≤≥⪯⪰  ∼≈⊂⊃≪≫≺≻
032  ←→↑↓↔↗↘≃  ⇐⇒⇑⇓⇔↖↙∝  ′∞∈∋△▽  ∀̸∃¬∅ℜℑ⊤⊥
064  ℵABCDEFG  HIJKLMNO  PQRSTUVW  XYZ∪∩⊎∧∨
096  ⊢⊣⌊⌋⌈⌉{}  ⟨⟩|∥↕⇕\≀  √⨿∇∫⊔⊓⊑⊒  §†‡¶♣♢♡♠

12. cmbx10

000  ΓΔΘΛΞΠΣΥ  ΦΨΩfffiflffiffl  ıȷ` ´ ˇ ˘  ¯  ˚  ¸ ßæœøÆŒØ
032 !”#$%&’  ()*+,-./  01234567  89:;¡=¿?
064  @ABCDEFG  HIJKLMNO  PQRSTUVW  XYZ[“]ˆ ˙
096  ‘abcdefg  hijklmno  pqrstuvw  xyz–—˝ ˜ ¨

13. cmbx7

000  ΓΔΘΛΞΠΣΥ  ΦΨΩfffiflffiffl  ıȷ` ´ ˇ ˘  ¯  ˚  ¸ ßæœøÆŒØ
032 !”#$%&’  ()*+,-./  01234567  89:;¡=¿?
064  @ABCDEFG  HIJKLMNO  PQRSTUVW  XYZ[“]ˆ ˙
096  ‘abcdefg  hijklmno  pqrstuvw  xyz–—˝ ˜ ¨

14. cmbx5

000  ΓΔΘΛΞΠΣΥ  ΦΨΩfffiflffiffl  ıȷ` ´ ˇ  ˘  ¯  ˚  ¸ ßæœøÆŒØ
032 !”#$%&’  ()*+,-./  01234567  89:;¡=¿?
064  @ABCDEFG  HIJKLMNO  PQRSTUVW  XYZ[“]ˆ ˙
096  ‘abcdefg  hijklmno  pqrstuvw  xyz–—˝ ˜ ¨

15. cmti10

000  ΓΔΘΛΞΠΣΥ  ΦΨΩfffiflffiffl  ıȷ` ´ ˇ ˘  ¯  ˚  ¸ ßæœøÆŒØ
032 !”#£%&’  ()*+,-./  01234567  89:;¡=¿?
064  @ABCDEFG  HIJKLMNO  PQRSTUVW  XYZ[“]ˆ ˙
096  ‘abcdefg  hijklmno  pqrstuvw  xyz–—˝ ˜ ¨

16. cmex10

000  ()[]⌊⌋⌈⌉  {}⟨⟩|∥/\  ()()[]⌊⌋  ⌈⌉{}⟨⟩/\

032  ()[]⌊⌋⌈⌉  {}⟨⟩/\/\

064  ⟨⟩⨆⨆  ∮∮⨀⨀⨁⨁⨂⨂  ∑∏∫⋃⋂⨄⋀⋁  ∑∏∫⋃⋂⨄⋀⋁

096  ∐∐  []⌊⌋⌈⌉{}  √√√√√

\sqrt{1 + \sqrt{1 + \sqrt{1 + \sqrt{1 + \sqrt{1 + \sqrt{1 + \sqrt{1 + \sqrt{1 + \sqrt{1 + \sqrt{1 + \sqrt{1 + x}}}}}}}}}}}